Yuri has a board that is 98 inches long. He wishes to cut the board into two pieces so that one piece will be 10 inches longer than the other.
Equation Used:
Length of piece 1:

Length of piece 2:

Answer by checkley71(8403)

X+(X+10)=98
X+X=10=98
2X=98-10
2X=88
X=88/2
X=44 INCHES FOR THE SHORTER PIECE
98-44=54 FOR THE LONGER PIECE.
PROOF
44+54=98 AND
54=44+10